# Measuring multipartite quantum correlations by thermodynamic work extraction

### v1

@@ D.3 Tightness of the lower and upper bounds
This inequality shows that we can achieve arbitrary accuracy $\varepsilon$ by increasing the locality $l$. In the derivation, we utilize the measurement protocol $\Lambda^{*}_{n:l}\in\mathcal{N}_{n:l}$, which achieves the minimum for $\min_{\Lambda\in\mathcal{N}_{n:l}}H_{\Lambda}(Y_{n:l}|Z_{n})$. With this notation, we can derive the inequality (14) as
@@
$\leq\frac{1}{l}2\ln D_{B}.$
Here, $H_{\Lambda}(y:z)$ represents the classical mutual information between measurement outcomes $y$ and $z$ under the measurement protocol $\Lambda$, and $I_{a_{1}:a_{2}}(\rho_{1,2})$ represents the quantum mutual information between party $a_{1}$ and $a_{2}$ of the state $\rho_{1,2}$. The inequality in the third line follows from the fact that quantum mutual information decreases monotonically under local CPTP map, which is the consequence of the monotonicity of quantum relative entropy Petz (2003). To show the inequality in the fifth line, we suppose that quantum state is described with the MPS with a fixed bond dimension $D_{B}$.
Additionally, we can derive the following inequalities, which hold in general, not only for MPS:
@@ Appendix E Efficient calculation method for one-way work deficit density $\delta^{\to}$
E.1 Lower bounds
We here introduce the lower bound for one-way work deficit density $L_{l}^{\to}$. The lower bound is defined as
@@ E.1 Lower bounds
where $\mathcal{N}_{n:l}^{\to}$ is the set of measurements whose POVM operators are represented as
@@
Here, the local measurement on party $a_{n+k+1}$ depends on the past measurement outcomes on parties $\{a_{i}\}_{i=1}^{n+k}$ (i.e., $Z_{n}$ and $Y_{n:k}$). Similarly to $L_{l}$, the optimization region of the POVM $\{E_{Z_{n}}\}$ can be reduced as shown in Fig. 12 for systems described with MPS. Furthermore, the continuity of the lower bound can be also proved as follows similarly to Eq. (55):
$|L_{l}^{\to}(\rho)-L_{l}^{\to}(\rho^{\prime})|\leq\tilde{\mathcal{O}}(\mu),$ (61)
for states $\rho$ and $\rho^{\prime}$ satisfying $\|\rho-\rho^{\prime}\|_{\rm tr}\leq\mu$. Since each term in the summation of Eq. (60), $\min_{\Lambda\in\mathcal{N}^{\to}_{n:l}}H_{\Lambda}(Y_{n:l}|Z_{n})$, can be calculated with the constant numerical cost, when the system is described with MPS, the bound $L_{l}^{\to}$ can be calculated with the cost proportional to the system size $N$. This lower bound $L_{l}^{\to}$ is derived as follows:
@@
for arbitrary natural numver $k$, which implies that the bound $L_{l}^{\to}$ is systematically improved by increasing the locality parameter $l$. The derivation is essentially the same as Eq. (58).
E.2 Upper bounds
We here introduce the upper bounds $U_{l}^{\to}$ for the one-way work deficit density. The bound $U_{l}^{\to}$ is defined as
Here, $\{\Pi_{y_{n+k+1}}^{Y_{n:k}}\}_{Y_{n:k+1}}$ represents the local projective measurement on $a_{n+k+1}$ under the condition that the measurement outcomes on $\{a_{i}\}_{i=n+1}^{n+k}$ are $Y_{n:k}$. Since each term in the summation of Eq. (63), $\min_{\Lambda\in\mathcal{M}_{n:l}^{\to}}H_{\Lambda}(Y_{n:l})$, can be calculated with the constant numerical cost, the bound $U_{l}^{\to}$ is calculated with the cost proportional to the system size $N$. This bound is derived as follows:
$=U_{l}^{\to}.$
Furthermore, we can show the inequality
$U_{l}^{\to}\geq U_{kl}^{\to},$ (64)
for arbitrary natural numver $k$, which implies that the bound $U_{l}^{\to}$ is systematically improved by increasing the locality parameter $l$. The derivation is essentially the same as Eq. (59).

### v2

@@ D.4 Tightness of the lower and upper bounds
This inequality shows that we can achieve arbitrary accuracy $\varepsilon$ by increasing the locality $l$. In the derivation, we utilize the measurement protocol $\Lambda^{*}_{n:l}\in\mathcal{N}_{n:l}$, which achieves the minimum for $\min_{\Lambda\in\mathcal{N}_{n:l}}H_{\Lambda}(Y_{n:l}|Z_{n})$. With this notation, we can derive the inequality (14) as follows:
@@
$\leq\frac{1}{l}4\ln D_{B}=\mathcal{O}\left(\frac{1}{l}\right).$
Here, $H_{\Lambda}(y:z)\equiv H_{\Lambda}(y)+H_{\Lambda}(z)-H_{\Lambda}(y,z)$ represents the classical mutual information between measurement outcomes $y$ and $z$ under the measurement protocol $\Lambda$, and $I_{a_{1}:a_{2}}(\rho)\equiv S(\rho_{1})+S(\rho_{2})-S(\rho_{1,2})$ represents the quantum mutual information between parties $a_{1}$ and $a_{2}$. The inequality in the third line follows from the fact that quantum mutual information decreases monotonically under local CPTP map, which is the consequence of the monotonicity of quantum relative entropy Petz (2003). In the derivation of the final line, we utilize that the system is described with an MPS with a fixed bond dimension $D_{B}$.
Additionally, we can derive the following inequalities, which are satisfied in general, not only in MPS:
@@ Appendix E Efficient calculation method for one-way work deficit density $\delta^{\to}$
E.1 Upper bounds
We here introduce the upper bound $U_{l}^{\to}$ for the one-way work deficit density $\delta^{\to}$. The bound $U_{l}^{\to}$ is defined as
Here, $\{\Pi_{y_{n+k+1}}^{Y_{n:k}}\}_{Y_{n:k+1}}$ represents the local projective measurement on $a_{n+k+1}$ under the condition that the measurement outcomes on $\{a_{i}\}_{i=n+1}^{n+k}$ are $Y_{n:k}$. Since each term in the summation of Eq. (63), $\min_{\Lambda\in\mathcal{M}_{n:l}^{\to}}H_{\Lambda}(Y_{n:l})$, can be calculated at a constant numerical cost independent of system size $N$, the bound $U_{l}^{\to}$ is calculated with the cost proportional to $N$.
The upper bound $U_{l}^{\to}$ is derived as follows:
$=U_{l}^{\to}.$
Furthermore, we can show the inequality
$U_{l}^{\to}\geq U_{kl}^{\to},$ (64)
for an arbitrary natural number $k$, which implies that the bound $U_{l}^{\to}$ is systematically improved by increasing the locality parameter $l$. The derivation of Eq. (66) is essentially the same as Eq. (62).
E.2 Lower bounds
We here introduce the lower bound $L_{l}^{\to}$ for one-way work deficit density $\delta^{\to}$. The lower bound is defined as
@@ E.2 Lower bounds
where $\mathcal{N}_{n:l}^{\to}$ is the set of measurements whose operators are represented as
@@
Here, the local measurement on party $a_{n+k+1}$ depends on the past measurement outcomes on parties $\{a_{i}\}_{i=1}^{n+k}$ (i.e., $Z_{n}$ and $Y_{n:k}$). Similarly to $L_{l}$, the optimization region of the POVM $\{E_{Z_{n}}\}$ can be reduced as shown in Fig. 12 for systems described with MPS. Furthermore, the continuity of the lower bound can also be proved, similarly to Eq. (60). Since each term in the summation of Eq. (65), $\min_{\Lambda\in\mathcal{N}^{\to}_{n:l}}H_{\Lambda}(Y_{n:l}|Z_{n})$, can be calculated with the constant numerical cost for systems described with MPS, the bound $L_{l}^{\to}$ can be calculated with the cost proportional to the system size $N$.
The lower bound $L_{l}^{\to}$ is derived as follows:
@@
for an arbitrary natural number $k$, which implies that the bound $L_{l}^{\to}$ is systematically improved by increasing the locality parameter $l$. The derivation of Eq. (66) is essentially the same as Eq. (61).